a body released from a great height falls freely towards earth.another body is released from the same height exactly one second later.then the seperation between two bodies after release of two second body is

Grade:11

2 Answers

Arun
6 years ago

Displacement of the first body after 3 seconds= S = ut + ½ g t^2  = 44.1m/s

Displacement of the first body after 2 seconds= S = ut + ut + ½ g t^2  = 19.6m/s

The separation between them=44.5 - 19.6 = 24.5m/s

Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Displacement of the first body after 3 seconds =
S = ut + ½ g t^2
= 44.1m/s
Displacement of the first body after 2 seconds=
S = ut + ½ g t^2
= 19.6m/s
The separation between them
= 44.5 - 19.6
= 24.5m/s

Thanks and Regards